OpenStudy (ria23):

Identify the location of the x-intercept(s), if any. Leave yhur answers in simplified radical form. y=x^2-x-12

OpenStudy (ria23):

x^2-x-12=0 Factor: (x-4)(x+3)=0 That's as far as I got... Now I'm stuck.

OpenStudy (anonymous):

Set each one equal to 0 and solve for x

OpenStudy (anonymous):

x-4 = 0 and x + 3 = 0

OpenStudy (ria23):

Oh... 4 and -3. c: Thank yhu! Again. ^.^

OpenStudy (anonymous):

yep and if they want points the y-values to each is just 0 so (4,0) and (-3,0)

OpenStudy (ria23):

Thank yhu C: